What is the net emf of different cell connected in parallel GIVEN THAT EVERY CELL HAS ZERO INTERNAL RESISTANCE

Grade:12

2 Answers

Arun
5 years ago

Let the equivalent emf be E

and the equivalent resistance be r

From parallel combination of resistors, we can deduce that r= r1r2/ (r1+r2)

Also, using KCL, I=I1+I2

From Ohm’s Law, we know that, V=IR

For first branch,

I1= E1/r1

Similarly, for second branch,

I2=E2/r2

Therefore, the Equivalent emf, E= Ir= (E1/r1 + E2/r2)* r1r1/(r1+r2) = (E1r2+E2r1)/ (r1+r2)

Manav Patel
5 years ago
But according to this formula if INTERNAL RESISTANCE = 0 then net emf will br infinity?...
I know this formula but is it true for ZERO INTERNAL RESISTANCE? .......